# Tight closure of products and F-rational singularities

### v3

@@ 1. Introduction
For many purposes, F-rational singularities can be regarded as the “correct” characteristic $p>0$ analogue of rational singularities in characteristic 00. It is known that F-rational singularities are normal, Cohen-Macaulay and pseudo-rational [Smi97]. Our main result is the following.
@@ Acknowledgements
We thank Craig Huneke and Pham Hung Quy for helpful discussions and Burt Totaro for providing feedback on an earlier version of this preprint. We also thank Holger Brenner, Craig Huneke, Vivek Mukundan and Jugal Verma for allowing us to include Theorem 3.10 in this paper. The first named author was partially supported by the MIUR Excellence Department Project CUP D33C23001110001, and by INdAM-GNSAGA. The second author was supported by the Ramon y Cajal Fellowship RYC2020-028976-I funded by MCIN/AEI/10.13039/501100011033 and by FSE “invest in your future” and projects PID2021-125052NA-I00 funded by MCIN/AEI/10.13039/501100011033 and EUR2023-143443 funded by MCIN/AEI/10.13039/501100011033 and the European Union NextGenerationEU/PRTR.
@@ 2. Background
Throughout this article, $(R,\mathfrak{m})$ denotes a local ring of prime characteristic $p>0$. Given a proper ideal $I\subseteq R$, and $q=p^{e}$, we let $I^{[q]}=(y^{q}\mid y\in I)$ be its Frobenius power. Note that $I^{[q]}=F^{e}(I)R$, where $F\colon R\to R$ is the Frobenius map, that is, the ring homomorphism raising an element to its $p$-th power.
@@
Let $(R,\mathfrak{m})$ be an F-injective equidimensional local ring which F-rational on the punctured spectrum. Then any parameter ideal $\mathfrak{q}$ is Frobenius closed and $\mathfrak{m}\mathfrak{q}^{*}\subseteq\mathfrak{q}$.
@@ 3. Tight closure of products of parameter ideals
If $R$ is F-rational there is nothing to prove, so assume that $R$ is not F-rational. By Corollary 2.9 we have that $\mathfrak{q}^{*}/\mathfrak{q}$ is an $R/\mathfrak{m}$-vector space. Let $u_{1},\ldots,u_{t}\in\mathfrak{q}^{*}$ be a basis modulo $\mathfrak{q}$. By the Main Theorem of [Quy18], we have that $\lambda((\mathfrak{q}^{[q]})^{*}/\mathfrak{q}^{[q]})$ does not depend on $q$. Since $u_{1}^{q},\ldots,u_{t}^{q}$ belong to $(\mathfrak{q}^{*})^{[q]}\subseteq(\mathfrak{q}^{[q]})^{*}$, and because $(\mathfrak{q}^{[q]})^{*}/\mathfrak{q}^{[q]}$ is an $R/\mathfrak{m}$-vector space of dimension $t=\ell(\mathfrak{q}^{*}/\mathfrak{q})$, to complete the proof it suffices to show that $u_{1}^{q},\ldots,u_{t}^{q}$ are linearly independent modulo $\mathfrak{q}^{[q]}$. To this end, assume that $\lambda_{1}u_{1}^{q}+\ldots+\lambda_{t}u_{t}^{q}\in\mathfrak{q}^{[q]}$. We want to show that $\lambda_{i}\in\mathfrak{m}$ for all $i$. Since the residue field is perfect, we can choose $\alpha_{1},\ldots,\alpha_{t}\in R$ such that $\alpha_{j}^{q}\equiv\lambda_{j}$ modulo $\mathfrak{m}$. By Corollary 2.9 we have that $\mathfrak{m}u_{j}^{q}\subseteq\mathfrak{m}(\mathfrak{q}^{[q]})^{*}\subseteq\mathfrak{q}^{[q]}$ for all $j$, and therefore we get that $(\alpha_{1}u_{1}+\ldots+\alpha_{t}u_{t})^{q}\equiv\lambda_{1}u_{1}^{q}+\ldots+\lambda_{t}u_{t}^{q}\equiv 0$ modulo $\mathfrak{q}^{[q]}$. Since $R$ is equidimensional and F-rational on the punctured spectrum, the local cohomology modules $H^{i}_{\mathfrak{m}}(R)$ have finite length for all $i\neq\dim(R)$. Since $R$ is F-injective we then get that $\alpha_{1}u_{1}+\ldots+\alpha_{t}u_{t}\in\mathfrak{q}^{F}$, and moreover $\mathfrak{q}^{F}=\mathfrak{q}$ by Corollary 2.9. It follows that $\alpha_{1},\ldots,\alpha_{t}\in\mathfrak{m}$. Finally, this implies that $\lambda_{1},\ldots,\lambda_{t}\in\mathfrak{m}$, and finishes the proof. ∎
@@
We now reduced to the case in which $R_{\mathfrak{p}}$ is an F-injective ring which is F-rational on the punctured spectrum and the parameter ideal $\mathfrak{a}=\mathfrak{q}R_{\mathfrak{p}}$ in $R_{\mathfrak{p}}$ and is such that $(\mathfrak{a}\mathfrak{a}^{[q]})^{*}=\mathfrak{a}^{*}(\mathfrak{a}^{[q]})^{*}$ for all $q=p^{e}$. Theorem 3.3 implies that $R_{\mathfrak{p}}$ is F-rational; a contradiction. ∎
Trying to find analogies between Theorem 3.9 and [Cut90, Theorem 1], one might wonder whether the condition $(IJ)^{*}=I^{*}J^{*}$ for all ideals $I,J\subseteq R$ can be equivalent to $R$ being F-rational. This is actually not the case: the following argument was shown to us in a private communication by Brenner, Huneke, Mukundan, and Verma.
@@
Let $(R,\mathfrak{m})$ be an F-rational local ring. The following are equivalent:
@@
The implications (1) $\Rightarrow$ (2) $\Rightarrow$ (3) are straightforward. Assume (3). In order to show that $R$ is weakly F-regular it suffices to prove that every $\mathfrak{m}$-primary ideal is tightly closed. Let $I$ be an $\mathfrak{m}$-primary ideal, we induct on $n=\mu(I)\geqslant d=\dim(R)$. If $n=d$ then $I$ is a parameter ideal, and hence it is tightly closed since $R$ is F-rational. Assume $n>d$. By prime avoidance, we can find generators $x_{1},\ldots,x_{n}$ of $I$ such that $J=(x_{1},\ldots,x_{d})$ is a parameter ideal, and $K=(x_{1},\ldots,x_{d-1},x_{d+1},\ldots,x_{n})$ is still $\mathfrak{m}$-primary. In particular, we have that $K^{*}=K$ by induction since $\mu(K)<n$. Let $z\in I^{*}$. There exists $c\in R^{\circ}$ such that $cz^{q}\in I^{[q]}$ for all $q\gg 0$. Multiplying by $x_{1}^{q}$ we get that $cx_{1}^{q}z^{q}\in x_{1}^{q}I^{[q]}\subseteq J^{[q]}K^{[q]}$. It follows that $x_{1}z\in(JK)^{*}=JK^{*}=JK$. Write $x_{1}z=\sum_{i=1}^{d}a_{i}x_{i}$ with $a_{i}\in K\subseteq I$, so that $(z-a_{1})x_{1}\in(x_{2},\ldots,x_{d})$. Since $x_{1},\ldots,x_{d}$ is a regular sequence we conclude that $z\in(a_{1},x_{2},\ldots,x_{d})\subseteq I$, as desired. ∎
@@ 4. Ending remarks and questions
Let $(R,\mathfrak{m})$ be a two-dimensional excellent F-injective analytically normal local ring. Is being F-rational equivalent to the condition $(I^{2})^{*}=(I^{*})^{2}$ for all $\mathfrak{m}$-primary ideals $I\subseteq R$? Or perhaps to $(\mathfrak{q}^{2})^{*}=(\mathfrak{q}^{*})^{2}$ for all ideals $\mathfrak{q}$ generated by a system of parameters?
Question 4.4.
Combining the Theorem of Huneke–Itoh [Hun87, Ito88] with the aforementioned result of Lipman–Teissier on the Briançon–Skoda Theorem, we know that a two-dimensional normal local ring is pseudo-rational if and only if $\overline{\mathfrak{q}^{2}}\subseteq\mathfrak{q}$ for every ideal $\mathfrak{q}$ generated by a system of parameters. In light of this equivalence, and keeping in mind the tight closure approach to the Briançon–Skoda Theorem [HH90, AH01], a natural question arises: is being F-rational equivalent to $(\mathfrak{q}^{2})^{*}\subseteq\mathfrak{q}$ for every ideal $\mathfrak{q}$ generated by a system of parameters?

### v4

@@ 1. Introduction
For many purposes, F-rational singularities can be regarded as the “correct” characteristic $p>0$ analogue of rational singularities in characteristic 0. It is known that F-rational singularities are normal, Cohen-Macaulay and pseudo-rational [Smi97]. Our main result is the following.
@@ Acknowledgements
We thank Craig Huneke, Pham Hung Quy and Florian Enescu for helpful discussions, and Burt Totaro for providing feedback on an earlier version of this preprint. We also thank the referee for helping us improve the quality of the exposition. We are grateful to Holger Brenner, Craig Huneke, Vivek Mukundan and Jugal Verma for allowing us to include Theorem 3.10 in this paper. The first named author was partially supported by the MIUR Excellence Department Project CUP D33C23001110001, and by INdAM-GNSAGA. The second author was supported by Spanish Ministry of Science, Innovation, and Universities under grants PCI2024-155055-2 funded by MICIU/AEI/10.13039/501100011033 and UE, RYC2020-028976-I funded by MICIU/AEI/10.13039/501100011033 and by FSE “invest in your future”, and EUR2023-143443 funded by MICIU/AEI/10.13039/501100011033 and the European Union NextGenerationEU/PRTR.
@@ 2. Background
Throughout this article, $(R,\mathfrak{m})$ denotes a local ring of prime characteristic $p>0$. Given a proper ideal $I\subseteq R$, and $q=p^{e}$, we let $I^{[q]}=(y^{q}\mid y\in I)$ be its Frobenius power. Note that $I^{[q]}=F^{e}(I)R$, where $F\colon R\to R$ is the Frobenius map, that is, the ring homomorphism raising an element to its $p$-th power. We will use $\ell_{R}(-)$ to denote the length of an $R$-module.
@@
Let $(R,\mathfrak{m})$ be an F-injective equidimensional local ring which is F-rational on the punctured spectrum. Then any parameter ideal $\mathfrak{q}$ is Frobenius closed and $\mathfrak{m}\mathfrak{q}^{*}\subseteq\mathfrak{q}$.
@@ 3. Tight closure of products of parameter ideals
If $R$ is F-rational there is nothing to prove, so assume that $R$ is not F-rational. By Corollary 2.9 we have that $\mathfrak{q}^{*}/\mathfrak{q}$ is an $R/\mathfrak{m}$-vector space. Let $u_{1},\ldots,u_{t}\in\mathfrak{q}^{*}$ be a basis modulo $\mathfrak{q}$. By the Main Theorem of [Quy18], we have that $\ell_{R}((\mathfrak{q}^{[q]})^{*}/\mathfrak{q}^{[q]})$ does not depend on $q$. Since $u_{1}^{q},\ldots,u_{t}^{q}$ belong to $(\mathfrak{q}^{*})^{[q]}\subseteq(\mathfrak{q}^{[q]})^{*}$, and because $(\mathfrak{q}^{[q]})^{*}/\mathfrak{q}^{[q]}$ is an $R/\mathfrak{m}$-vector space of dimension $t=\ell_{R}(\mathfrak{q}^{*}/\mathfrak{q})$, to complete the proof it suffices to show that $u_{1}^{q},\ldots,u_{t}^{q}$ are linearly independent modulo $\mathfrak{q}^{[q]}$. To this end, assume that $\lambda_{1}u_{1}^{q}+\ldots+\lambda_{t}u_{t}^{q}\in\mathfrak{q}^{[q]}$. We want to show that $\lambda_{i}\in\mathfrak{m}$ for all $i$. Since the residue field is perfect, we can choose $\alpha_{1},\ldots,\alpha_{t}\in R$ such that $\alpha_{j}^{q}\equiv\lambda_{j}$ modulo $\mathfrak{m}$. By Corollary 2.9 we have that $\mathfrak{m}u_{j}^{q}\subseteq\mathfrak{m}(\mathfrak{q}^{[q]})^{*}\subseteq\mathfrak{q}^{[q]}$ for all $j$, and therefore we get that $(\alpha_{1}u_{1}+\ldots+\alpha_{t}u_{t})^{q}\equiv\lambda_{1}u_{1}^{q}+\ldots+\lambda_{t}u_{t}^{q}\equiv 0$ modulo $\mathfrak{q}^{[q]}$, that is, $\alpha_{1}u_{1}+\ldots+\alpha_{t}u_{t}\in\mathfrak{q}^{F}$. By Corollary 2.9 we conclude that $\alpha_{1}u_{1}+\ldots+\alpha_{t}u_{t}\in\mathfrak{q}$, and it follows from our choice of $u_{1},\ldots,u_{t}$ that $\alpha_{1},\ldots,\alpha_{t}\in\mathfrak{m}$. Finally, this implies that $\lambda_{1},\ldots,\lambda_{t}\in\mathfrak{m}$, and finishes the proof. ∎
@@
We now reduced to the case in which $R_{\mathfrak{p}}$ is an F-injective ring which is F-rational on the punctured spectrum and the parameter ideal $\mathfrak{a}=\mathfrak{q}R_{\mathfrak{p}}$ in $R_{\mathfrak{p}}$ and is such that $(\mathfrak{a}\mathfrak{a}^{[q]})^{*}=\mathfrak{a}^{*}(\mathfrak{a}^{[q]})^{*}$ for all $q=p^{e}$. Theorem 3.3 implies that $R_{\mathfrak{p}}$ is F-rational, a contradiction. ∎
In trying to find analogies between Theorem 3.9 and [Cut90, Theorem 1], one might wonder whether the condition $(IJ)^{*}=I^{*}J^{*}$ for all ideals $I,J\subseteq R$ may be equivalent to $R$ being F-rational. This is actually not the case: the following argument was shown to us in a private communication by Brenner, Huneke, Mukundan, and Verma.
@@
Let $(R,\mathfrak{m})$ be an F-rational local ring of dimension $d$. The following are equivalent:
@@
The implications (1) $\Rightarrow$ (2) $\Rightarrow$ (3) are straightforward. Assume (3). In order to show that $R$ is weakly F-regular it suffices to prove that every $\mathfrak{m}$-primary ideal is tightly closed. Let $I$ be an $\mathfrak{m}$-primary ideal, we induct on $n=\mu(I)\geqslant d=\dim(R)$. If $n=d$ then $I$ is a parameter ideal, and hence it is tightly closed since $R$ is F-rational. Assume $n>d$. We claim that, we can find minimal generators $x_{1},\ldots,x_{n}$ of $I$ such that $J=(x_{1},\ldots,x_{d})$ is a parameter ideal, and $K=(x_{1},\ldots,x_{d-1},x_{d+1},\ldots,x_{n})$ is still $\mathfrak{m}$-primary. In fact, choose any parameter ideal $J=(x_{1},\ldots,x_{d})$ inside $I$. Observe that $I\not\subseteq J+\mathfrak{m}I$, othwerwise we would have $I=J$ by Nakayama’s lemma, and thus $n=d$. Moreover, for any minimal prime $\mathfrak{p}$ of $(x_{1},\ldots,x_{d-1})$, we have $I\not\subseteq\mathfrak{p}$, since otherwise $\operatorname{ht}(I)\leqslant\operatorname{ht}(\mathfrak{p})=d-1$. It follows by prime avoidance that we can find $x_{d+1}\in I\smallsetminus(J+\mathfrak{m}I)$ such that $x_{d+1}\notin\mathfrak{p}$ for all minimal primes $\mathfrak{p}$ of $(x_{1},\ldots,x_{d-1})$. The latter condition ensures that $(x_{1},\ldots,x_{d-1},x_{d+1})$ is $\mathfrak{m}$-primary. Moreover, our choice of $x_{d+1}$ guarantees that the residue class $\overline{x_{d+1}}\in I/\mathfrak{m}I$ does not belong to the $R/\mathfrak{m}$-vector subspace spanned by $\overline{x_{1}},\ldots,\overline{x_{d}}$. By Nakayama’s lemma this implies that $x_{1},\ldots,x_{d},x_{d+1}$ are part of a minimal generating set for $I$. Now complete $\overline{x_{1}},\ldots,\overline{x_{d}},\overline{x_{d+1}}$ to a $R/\mathfrak{m}$-basis of $I/\mathfrak{m}I$ with elements $\overline{x_{d+2}},\ldots,\overline{x_{n}}$, and the elements $x_{1},\ldots,x_{d},x_{d+1},\ldots,x_{n}$ satisfy the desired properties.
Now note that $\mu(K)<n$, so by induction we have that $K^{*}=K$. Let $z\in I^{*}$. There exists $c\in R^{\circ}$ such that $cz^{q}\in I^{[q]}$ for all $q\gg 0$. Multiplying by $x_{1}^{q}$ we get that $cx_{1}^{q}z^{q}\in x_{1}^{q}I^{[q]}\subseteq J^{[q]}K^{[q]}$. It follows that $x_{1}z\in(JK)^{*}=JK^{*}=JK$. Write $x_{1}z=\sum_{i=1}^{d}a_{i}x_{i}$ with $a_{i}\in K\subseteq I$, so that $(z-a_{1})x_{1}\in(x_{2},\ldots,x_{d})$. Since $x_{1},\ldots,x_{d}$ is a regular sequence, we conclude that $z\in(a_{1},x_{2},\ldots,x_{d})\subseteq I$, as desired. ∎
@@ 4. Ending remarks and questions
Let $(R,\mathfrak{m})$ be a two-dimensional excellent F-injective analytically normal local ring. Is being F-rational equivalent to the condition $(I^{2})^{*}=(I^{*})^{2}$ for all $\mathfrak{m}$-primary ideals $I\subseteq R$?
Remark 4.4.
We point out that the condition that $(\mathfrak{q}^{2})^{*}=(\mathfrak{q}^{*})^{2}$ for all ideals $\mathfrak{q}\subseteq R$ generated by a full system of parameters does not imply that $R$ is F-rational. For example, take $R$ to be a two-dimensional Gorenstein pseudo-rational F-injective local ring. We point out that such a ring needs not be F-rational; for instance, take $R=\mathbb{F}_{2}\llbracket{x,y,z}\rrbracket/(z^{2}+x^{2}y+xy^{2}+xyz)$. Let $\mathfrak{q}=(f,g)$ be a parameter ideal of $R$. Since the test ideal $\tau(R)$ equals $\mathfrak{m}$ we have $I^{*}\subseteq I:\mathfrak{m}$ for any $I$. Furthermore, one always has $(\mathfrak{q}^{*})^{2}\subseteq(\mathfrak{q}^{2})^{*}$, so
$(\mathfrak{q}:\mathfrak{m})^{2}=(\mathfrak{q}^{*})^{2}\subseteq(\mathfrak{q}^{2})^{*}\subseteq\mathfrak{q}^{2}:\mathfrak{m},$
Now, if we let $\mathfrak{q}:\mathfrak{m}=(f,g,u)$, then observe that $(f^{2},g):\mathfrak{m}=(f^{2},g,uf)$ and $(f,g^{2}):\mathfrak{m}=(f,g^{2},ug)$. From the injective map $0\to R/\mathfrak{q}^{2}\to R/(f^{2},g)\oplus R/(f,g^{2})$, one gets that the socle of $R/\mathfrak{q}^{2}$ is contained in the socle of $R/(f^{2},g)\oplus R/(f,g^{2})$, and this implies that $\mathfrak{q}^{2}:\mathfrak{m}\subseteq(\mathfrak{q}^{2},uf,ug)$. Thus $\mathfrak{q}^{2}:\mathfrak{m}\subseteq(\mathfrak{q}:\mathfrak{m})^{2}$, and we have equality $(\mathfrak{q}^{2})^{*}=(\mathfrak{q}^{*})^{2}$.
On a positive side, a similar argument shows that if $R$ is a Cohen-Macaulay then $\mathfrak{q}^{2}$ is tightly closed if and only if $\mathfrak{q}$ is tightly closed, i.e., $R$ is F-rational.
We now make another remark: combining the Theorem of Huneke–Itoh [Hun87, Ito88] with the aforementioned result of Lipman–Teissier on the Briançon–Skoda Theorem, we know that a two-dimensional normal local ring is pseudo-rational if and only if $\overline{\mathfrak{q}^{2}}\subseteq\mathfrak{q}$ for every ideal $\mathfrak{q}$ generated by a system of parameters. In light of this equivalence, and keeping in mind the tight closure approach to the Briançon–Skoda Theorem [HH90, AH01], another natural question arises: is being F-rational equivalent to $(\mathfrak{q}^{2})^{*}\subseteq\mathfrak{q}$ for every ideal $\mathfrak{q}$ generated by a system of parameters? However, this is false because one can take any pseudo-rational but not F-rational surface singularity, and exploit the containment $(\mathfrak{q}^{2})^{*}\subseteq\overline{\mathfrak{q}^{2}}$. For a concrete example, take $R=\mathbb{F}_{2}\llbracket{x,y,z}\rrbracket/(z^{2}+x^{2}y+xy^{2}+xyz)$ as above.